# LLMs and the Human Condition

### v3

@@ Abstract
Theory based AI research has had a hard time recently and the aim here is to propose a model of what LLMs are actually doing when they impress us with their language skills. The model integrates three established theories of human decision-making from philosophy, sociology, and computer science. The paper starts with the collective understanding of reasoning from the early days of AI research - primarily because that model is how we humans think we think, and is the most accessible. It then describes what is commonly thought of as "reactive systems" which is the position taken by many philosophers and indeed many contemporary AI researchers. The third component to the proposed model is from sociology and, although not flattering to our modern ego, provides an explanation to a puzzle that for many years has occupied those of us working on conversational user interfaces.
@@ 2 Machines
Computers are good at this kind of syntactic symbol manipulation, and between that and Schank’s observation that it is hard to think of a thousand things you know a hundred things about, this kind of reasoning formed the basis of thinking about computers and meaning. Today a spreadsheet is a great way to automate reasoning about some things. It turns out there is a wide range of tasks that fail when handled this way with a computer. Counting apples in a box might seem easy, but a box might have 15 apples in it according to Waitrose, 25 according to Aldi, and a horse would eat the lot no matter how many worms and rotten ones there are. Often the issue is speed, so if the idea is to map a room before you vacuum it, things that move like cats and people tend to re-set the model. In the early days of robot football the machines would wait for the ball to stop before deciding what to do. Air traffic control has a similar problem. At the time of writing aircraft have been grounded while air traffic controllers sort out why an aircraft, identified by flight number, occurred in two locations. There is something wrong with the system’s representation of the state of the world.
@@
The feeling in the late 1980s was that no matter how powerful computers might become in the future, simply getting the data for symbolic representations was going to be problematic. There were projects set up to encode all the data in the world, the CYC project Lenat et al. (1986) being the most explicit, but the scruffies were making progress and perhaps there was a better way. Rather than using sense-data to form a representation of the world, then reasoning about it before acting, functional systems could be built by connecting sensing to acting. Roomba vacuum cleaners do a random walk of a room to clean it and avoid chairs (and cats) by touching them and turning away. Since 1990 there has been a glacial shift in the computer science collective understanding of the problem and today there is wide spread acknowledgement Wooldridge (2023) that the nature of meaning is inseparably linked to our embodiment in the world. How this pans out for machines is still not clear and Brooks’ robots are often dismissed as being just “insect level intelligence” with robot developers continuing to add symbolic representations in order to support higher level intelligence. Philosophers on the other hand continue to push the limits of what higher level behaviours can be explained in terms of reactive systems Ward et al. (2017); Gallagher (2020); Bar-On (2021). The notion of a “reactive system” describes a mechanism used for AI problems, but it is not an interesting “algorithms and data-structures” solution. The ELIZA mechanism takes a set of if-then rules to map input to output, and applied to the DOCTOR script was found to be quite engaging. That same mechanism is used today for a myriad of chatbots but at the time and for many years after, AI researchers would consider it just a trick. From a philosopher’s perspective, the system is more interesting when considered as just a part of the agent-environment system. ELIZA worked in a particular environment in exactly the same way as Roombas work in a particular environment and, the claim is, in exactly the same way we humans make most of our decisions about what to do next.
@@ 3 Humans
We humans are not entirely reactive agents because we can, on occasion, think about things like whether I can afford that whiskey using an expression such as “£38 - £44 = …” but we also have another trick. We humans modify the environment to fit with us. In the cooking gnocchi for lunch scenario, none of the things I interact with are naturally occurring – someone made them. What is more, these things were put where I would find them. I might use insect level intelligence to eat lunch, but surely others must have had thoughts (with symbols) in order to create our benign environment. However it seems safe to assume other creatures do not.
@@
We do have the ability to do rational choice, and we do it with symbolic representations of things in the world. But following the strong version of the linguistic relativism argument, l’sign is arbitare Culler (1976) and what we reason about is culturally constructed. But the things all humans want to talk about are within contingencies, contingencies that are shared across all cultures. The suggestion is that we animals behave in accordance with both the reactive framework in which the environment “causes” action, and a framework of practices which evolved through evolutionary pressures. The unique human skill is the practice of “thingifying” the stuff around us, giving it a label, and then reasoning symbolically with the label. Language is both a window on how that is done, and a mechanism for passing it on.
@@ 4 A ’pragmatics first’ model of NLU
Whereas sociologists are interested at decision making at the macro level, the interest here is the mechanism for decision making at the agent level. Our aim is to find a plausible model of what goes on in heads. The position taken here is that language is indeed a window on the mind, but only the conscious part. We catch balls, and we catch a bus. We do not talk about the mechanism of eye-hand coordination directly, and we don’t “see” the intricate workings of the queue at the bus stop. For those of us interested in machines using language, we need to stop thinking about the meaning of words, and instead focus on the practices in which language is used to do things. Yes we do things with words Austin (1955), but the doing is embedded in practices.
@@
Of course in the example this does not happen. The child’s environment (i.e. Mother) does something different to what is expected at step two. In the child’s repertoire of practices however, mother’s utterance is a fit with the practice of checking homework. The child then has a decision to make: continue with the practice of getting something to eat, or to switch the PiP to checking homework. It is here that the child gets to do “purposive human action” – to reason about what to do next, The thing reasoned about is not his choice of words however, but about his choice of practice to perform. This level of decision is not linguistic - it is not about syntax and semantics, but about all sorts of non-linguistic but very pragmatic factors such as the power relation between mother and child.
In making the decision, the child can predict what will happen next by remembering how each practice played out in the past. Presented with line 11, (“I’ll let you have something later.”) the child knows what is coming next if he doesn’t stop niggling. We do not need to reason about other minds - the child does not need to reason about why mother is talking of homework - the child merely has to identify the practice in play and act out the appropriate role. Recognising the PiP is no more complex than recognising the words. Negotiating the PiP is where the interesting and I propose “conscious” decisions are being made.
@@ 7 Conclusion
Although we can and do think about what other people are thinking, we generally don’t bother. Much of what we do can be attributed to either reacting to stuff around us, or to the adoption of a role in a known practice. For social animals the practices are largely shared, and visible. That is, others can see what an individual is doing, and because the practice is shared, they know what the individual will do next. For humans, and the higher animals at least, practices can be learnt. The cat learns about the getting up in the morning routine; knows where the litter tray is and so on. People - language users - use language as part of practices, but language also empowers new practices. The symbolic nature of language reifies stuff into things so we can think about all cats when our experience is of just a few. We share these symbols - the sign and the signified - to invent (the practice of) counting and holding committee meetings, and we develop theories about "things" and how they interact. We can think about fairies, God, hermeneutics, and other people’s intentions.

### v4

@@ Abstract
Theory based AI research has had a hard time recently and the aim here is to propose a model of what LLMs are actually doing when they impress us with their language skills. The model integrates three established theories of human decision-making from philosophy, sociology, and computer science. The paper starts with the collective understanding of reasoning from the early days of AI research - primarily because that model is how we humans think we think. It then describes what is commonly thought of as "reactive systems" which is the position taken by many philosophers and indeed many contemporary AI researchers. The third component to the proposed model is from sociology and extends the reactive approach to communities as “agents” and devolves human intelligence to primarily an ecology of practices. Although not flattering to our modern ego the approach provides an explanation to a puzzle that for many years has occupied those of us working on conversational user interfaces.
@@ 2 Machines
Computers are good at this kind of syntactic symbol manipulation, and between that and Schank’s observation that it is hard to think of a thousand things you know a hundred things about, this kind of reasoning formed the basis of thinking about computers and meaning. Today a spreadsheet is a great way to automate reasoning about some things but it turns out there is a wide range of tasks that fail when handled this way. Counting apples in a box might seem easy, but a box might have 15 apples in it according to Waitrose, 25 according to Aldi, and a horse would eat the lot no matter how many are rotten. Modelling the world with referential symbols is hard because the referent can change. If the idea is to map a room before you vacuum it, cats and people tend move and re-set the model. In the early days of robot football the machines would wait for the ball to stop before deciding what to do. Air traffic control has a similar problem. At the time of writing aircraft have been grounded while air traffic controllers sort out why an aircraft, identified by flight number, occurred in two locations. There is something wrong with the system’s representation of the state of the world.
@@
The feeling in the late 1980s was that no matter how powerful computers might become in the future, simply getting the data for symbolic representations was going to be problematic. There were projects set up to encode all the data in the world, the CYC project Lenat et al. (1986) being the most explicit, but the scruffies were making progress and perhaps there was a better way. Rather than using sense-data to form a representation of the world, then reasoning about it before acting, functional systems could be built by connecting sensing to acting. Roomba vacuum cleaners do a random walk of a room to clean it and avoid chairs (and cats) by touching them and turning away. Since 1990 there has been a glacial shift in the computer science collective understanding of the problem and today there is wide spread acknowledgement Wooldridge (2023) that the nature of meaning is inseparably linked to our embodiment in the world.
How this pans out for machines is still not clear and Brooks’ robots are often dismissed as being just “insect level intelligence” with robot developers continuing to add symbolic representations in order to support higher level intelligence. Philosophers on the other hand continue to push the limits of what higher level behaviours can be explained in terms of reactive systems Hendriks-Jansen (1996); Ward et al. (2017); Gallagher (2020); Bar-On (2021). The notion of a “reactive system” describes a mechanism used for AI problems, but it is not an interesting “algorithms and data-structures” solution. The ELIZA mechanism takes a set of if-then rules to map input to output, and applied to the DOCTOR script was found to be quite engaging. That same mechanism is used today for a myriad of chatbots but at the time and for many years after, AI researchers would consider it just a trick.
From a philosopher’s perspective, the system is more interesting when considered as just a part of the agent-environment system. ELIZA worked in a particular environment in exactly the same way as Roombas work in a particular environment and, the claim is, in exactly the same way we humans make most of our decisions about what to do next.
@@ 3 Humans
We humans are not entirely reactive agents because we can, on occasion, think about things like whether I can afford that whiskey using an expression such as “£38 - £44 = …” but we also have another trick. We humans modify the environment to fit with us. In the cooking gnocchi for lunch scenario, none of the things I interact with are naturally occurring – someone made them. What is more, these things were put where I would find them. I might use insect level intelligence to eat lunch, but surely others must have had thoughts (with symbols) in order to create our benign environment.
@@
We do have the ability to do rational choice, and we do it with symbolic representations of things in the world, but following the strong version of the linguistic relativism argument, l’sign is arbitare Culler (1976) and what we reason about is culturally constructed. The reason human languages “carve reality” Lakoff (1987) in largely the same way is not because the world determines it, but because human interests are largely shared. The suggestion is that we animals behave in accordance with both the reactive framework in which the environment “causes” action, and a framework of practices which evolved through evolutionary pressures on our societies. The unique human skill is the practice of “thingifying” the stuff around us, giving it a label, and then reasoning symbolically with the label. Language is both a window on how that is done, and a mechanism for passing it on.
@@ 4 A ’pragmatics first’ model of NLU
Whereas sociologists are interested at decision making at the macro level, the interest here is the mechanism for decision making at the agent level. The aim is a model of what goes on in heads. Language is indeed a window on the mind, but only the conscious part. We catch balls, and we catch a bus. We do not talk about the mechanism of eye-hand coordination directly, and we don’t “see” the intricate workings of the queue at the bus stop. For those of us interested in machines using language, we need to stop thinking about the meaning of words, and instead focus on the practices in which language is used to do things. Yes we do things with words Austin (1955), but the doing is embedded in practices.
@@
Of course in the example this does not happen. The child’s environment (i.e. Mother) does something different to what is expected at step two. In the child’s repertoire of practices however, mother’s utterance is a fit with the practice of checking homework. The child then has a decision to make: continue with the practice of getting something to eat, or to switch the PiP to checking homework. It is here that the child gets to do “purposive human action” – to reason about what to do next, The thing reasoned about is not his choice of words however, but about his choice of practice to perform. This level of decision is not linguistic - it is not about syntax and semantics, but about all sorts of non-linguistic but very pragmatic factors such as the power relation between mother and child. In making the decision, the child can predict what will happen next by remembering how each practice played out in the past. Presented with line 11, (“I’ll let you have something later.”) the child knows what is coming next if he doesn’t stop niggling. The child does not need to reason about minds – about why mother is talking of homework – the child merely has to identify the practice in play and act out the appropriate role. Recognising the PiP is no more complex than recognising the words. Negotiating the PiP is where the interesting and I propose “conscious” decisions are being made.
@@ 6 What LLMs are actually doing
7 A human super power
Although we can and do think about what other people are thinking, we generally don’t bother. Much of what we do can be attributed to either reacting to stuff around us, or to the adoption of a role in a known practice. For social animals the practices are largely shared, and visible. That is, others can see what an individual is doing, and because the practice is shared, they know what the individual will do next. A sheep dog does not need to model the mind of a sheep – the sheep’s observable actions are simply a recognisable part of a practice that sheep perform (Figure 2).
Figure 2: Zak does’t read the Ewe’s mind, but instead sees a nasty head but unfolding.
For humans, and the higher animals at least, practices can be learnt. The cat learns about the getting up in the morning routine; knows where the litter tray is and so on. People - language users - use language as part of practices, but language also empowers new practices. The symbolic nature of language reifies stuff into things so we can think about all cats when our experience is of just a few. We share these symbols - the sign and the signified - to invent (the practice of) counting and holding committee meetings, and we develop theories about "things" and how they interact. We can think about fairies, God, hermeneutics, and other people’s intentions. The proposal is that it is generally only humans that do that. As language users we inherit a comprehensive set of signs and signified with which to think. It may be the case that other animals have the ability to do that as individuals, but language provides the toolbox which is handed down to the next generation along with the practices for using it.